The human body is approximately 60% of water, and one-half of the mass of the dry portion is protein. calculate the mass of protein in a 140. lb person.
1
Expert's answer
2019-08-27T04:43:49-0400

Percentage of dry portion in human body "=100-60=40\%"

percentage of protein in human body="\frac{40}2=20\%"

Mass of human body="140lb"

Mass of protein in human body"=\frac{20}{100}\times140=28lb"

now,

"1lb=0.45359237 kg"

"28lb=28\times0.45359237 =12.7kg"

So, mass of protein in 140"lb" person "=28lb=12.7kg"